OpenStudy (anonymous):

simplify cot^3(x)tan^4(x)cos(x)

OpenStudy (anonymous):

Is it sin(x)?

OpenStudy (kc_kennylau):

Convert everything in terms of sin(x) and cos(x)

OpenStudy (anonymous):

cot is the multiplicative inverse of tan and tan is sin/cos so eliassaab, yes

OpenStudy (kc_kennylau):

Hint: \(\tan x=\dfrac{\sin x}{\cos x}\)

OpenStudy (anonymous):

thanks :)